how to solve for this equation
x=9-2y
2x-y=3

Answer by kiru_khandelwal(79)

x = 9-2y ....(1)
2x - y = 3....(2)
Substituting the value of x from (1) in (2)
=> 2(9-2y) - y = 3
=> 18 -4y -y = 3
=> 18 -5y = 3
=> 5y = 15
=> y = 3
So x = 9-2*3 =3